OpenStudy (anonymous):

Solve the logarithmic equation for . Please round the answer to four decimal places, if necessary. ln(x-4)+ln(x+5)=1

OpenStudy (anonymous):

Reason I'm confused is because unlike log, in ln the exponential form is e^y=x and I'm having trouble solving the rest

OpenStudy (amistre64):

ln(a) + ln(b) = ln(ab)

OpenStudy (anonymous):

That's x^2+x-20?

OpenStudy (amistre64):

yep. now, ln(0) = 1 because e^0 = 1

myininaya (myininaya):

and also any number to the 0 power except 0 is 1

OpenStudy (amistre64):

your equation is really: x^2 +x - 20 = 0

myininaya (myininaya):

wait!

OpenStudy (amistre64):

or (x-4) (x+5) = 0

myininaya (myininaya):

e^1=e

myininaya (myininaya):

we have (x-4)(x+5)=e

OpenStudy (amistre64):

lol..... maybe it does, and maybe it doesnt :)

OpenStudy (anonymous):

yes myininaya

myininaya (myininaya):

if we do it like amistre we would make the problem much easier but also kindof wrong

myininaya (myininaya):

and kindof wrong is just wrong

OpenStudy (anonymous):

now im confused >_<

OpenStudy (amistre64):

horseshoe and handgranades me friend :)

myininaya (myininaya):

so we have x^2-9x+20+e=0

myininaya (myininaya):

oops -e

OpenStudy (anonymous):

wouldnt it be -e?

myininaya (myininaya):

Yes i blame amistre

myininaya (myininaya):

lol

OpenStudy (anonymous):

bring over 1 to the left at the start...before changinglg form...

OpenStudy (amistre64):

...its my liver pills...blame the liver pills :)

OpenStudy (anonymous):

so whats the answer?

myininaya (myininaya):

use quadractic formula x=(9@sqrt(81-4(1)(20-e))/2 where @ means + or -, but make sure your answer is positive because you cannot do ln(negative)

OpenStudy (anonymous):

sorry...doesn't make a difference what I said

OpenStudy (anonymous):

Quadratic formula the way to go

OpenStudy (anonymous):

this is what i put on the formula -1+squareroot(1^2 -4(1)(20-e))))/2

myininaya (myininaya):

no

myininaya (myininaya):

a=1 n=-9 c=20-e correct?

OpenStudy (anonymous):

c=-20-e

myininaya (myininaya):

b=-9*

OpenStudy (anonymous):

b=1

OpenStudy (anonymous):

no b =1

myininaya (myininaya):

dont listen to me

myininaya (myininaya):

im doing two things at once sorry

OpenStudy (anonymous):

my calculator wont work

myininaya (myininaya):

(-1+sqrt(1-4(-20-e))/2=4.2925 and (-1-sqrt(1-4(-20-e))/2=-5.2925 but we can't use -5.2925 since it is negative and the natural log of a negative number doesnt exist

OpenStudy (amistre64):

x can be negative if it is not the only thing controling the value of the log. the thing is: x-4 needs to be greater than 0 and, x+5 needs to be greater than 0. x >4 serves both cases.

myininaya (myininaya):

correct!

myininaya (myininaya):

that negative number gives you another negative number when you plug it into x-4